Operations Report – May 14th

Crew 194 Operations Report 14052018

SOL: 1

Name of person filing report: Dana

Non-nominal systems: Power/ Generator

Notes on non-nominal system:

functional but does not charge batteries

Solar—insufficient to power hab but does appear connected to batteries

Diesel – 75%

Gasoline (5 Gallon containers for ATV) – 15 gallons

Water (trailer) – 920 gallons

Water (static) – 55 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: nominal

ATV’s Used: 4

Oil Checked: yes

ATV # Fuel Used Gals: 7 gallons

Tires Status: good

# Hours Used Day: 2

Phobos used: yes

Hours: 2

Beginning charge: 100%

Ending charge: 95%

Currently charging: yes

Deimos used: No


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No


Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: No


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: No


Beginning charge:

Ending charge:

Currently charging:

HabCar used? No

Notes and Comments:

Summary of operations activities:

Summary of GreenHab activities: harvested Tomatoes and basil

Summary of ScienceDome activities: worked on construction of simulated Rocket

Summary of healthy and safety issues: none

Questions and Concerns to Mission Support: Power system remains off nominal. We are able to maintain electricity by using the diesel generator but the governor electrical system continues to not function.

Journalist Report – May 14th

Monday, May 14, 2018

Crew 194

Sol 1 Field Report

Morning lectures began early with details over space suits, hypoxia, toxicology. EVA immediately after with two teams successfully establishing comms relay further south of the Hab. Successful medical evac and treatment of injury due to fall. Afternoon consisted of rest, reading, Hab maintenance, and tidying. Crew all well.

Dean Aubin

Sol 1 Summary – May 14th

Crew 194 Daily Summary Report 14.05.2018

MDRS Daily Summary Report for sol 1

Summary Title

First EVA

Mission Status:


Sol Activity Summary:

Today was our first day in simulation. We learned about atmospheric composition and toxicology on spacecraft and habitats. We also discussed space suit design and had our first EVA.

Look Ahead Plan: We will continue our learning with discussion of radiation and celestial navigation.

Anomalies in work:

The Hab power system continue to be unreliable. We are on solar currently, but the system cuts out without warning. In addition, it does not appear to be sufficient to both power the Hab and recharge the batteries. The generator is able to power the Hab, but not to charge the batteries.

Weather:Sunny, Warm

Crew Physical Status:The crew is doing well. No concerns.

EVA: We had our first EVA today, which was exciting for the entire crew.

Reports to be file:

Operations Report

Support Requested:

Power System as above

Copyright © The Mars Society. All rights reserved. | Main Site